OpenStudy (anonymous):

A) estimate 51% of 162 explain how you got it B) use mental math to calculate 51% of 162 explain how u got it

OpenStudy (anonymous):

@DERRICKCURRY

OpenStudy (anonymous):

i am pretty sure it is A

OpenStudy (anonymous):

and do u have to answer the question too or whata

OpenStudy (anonymous):

no there both question

OpenStudy (anonymous):

u have to answer both of them

OpenStudy (anonymous):

To find a percentage: \[1\%=\frac{162}{100}=1.6.\]Now multiply it with 51. \[1.6*51=?\] NOTE: This is to calculate it the exact way. Since you need to find an estimate it would make sense just to take 50% which is half of 162. \[162*0.5=?\]

OpenStudy (anonymous):

This would be the answer to the first quest.

OpenStudy (anonymous):

I meant to write 1.62, and not 1.6

OpenStudy (welshfella):

50% of 162 = 1/2 of 162 = 81 1% of 162 = 1.6